OpenStudy (anonymous):

The size of a bicycle is determined by the diameter of the wheel. You want a 24-in bicycle for Christmas. What is the circumference of one of the wheels? A 75.4 in B 75.3 in C 37.7 in D 37.6 in

OpenStudy (anonymous):

@jdoe0001 do you think you can help with this one?

OpenStudy (jdoe0001):

"You want a 24-in bicycle for Christmas. " ^ diameter what would be the radius?

OpenStudy (anonymous):

the radius is 12 inches i think

OpenStudy (jdoe0001):

\(\bf circumference=2\pi {\color{red}{ r}}\qquad {\color{red}{ r}}=12\implies circumference=2\pi \cdot {\color{red}{ 12}} \)

OpenStudy (anonymous):

so the answer is a

OpenStudy (anonymous):

thanks for your help @jdoe0001 :)

OpenStudy (jdoe0001):

yw